OpenStudy (4meisu):

Rational Expressions: (1/2x) + (4x/x^2 - 1) - (2/x+1)

OpenStudy (anonymous):

you need to find the lcm of 2x, (x^2-1) and x+ 1

OpenStudy (anonymous):

first factor x^2-1 can you do this?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

i'm assuming you want to simplify this expression right?

OpenStudy (4meisu):

yes i do

OpenStudy (4meisu):

so the factor of x^2-1 will be (x-1) (x) ?

OpenStudy (anonymous):

x^2 - 1 this is 'difference' of two squares' = (x+1)(x-1) and the lcm = 2x(x+1)(x-1)

OpenStudy (anonymous):

now we multiply each of three expressions by lcm and obtain: (x+1)(x-1) +4x*2x - 2(x-1) ------------------------ 2x(x^2-1) whats left to do is simplify the top of the fraction

OpenStudy (4meisu):

thanks, my teacher didn't explain this concept well at all...

OpenStudy (anonymous):

sorry i made a slight error correct is (x+1)(x-1) +4x*2x - 4x(x-1) ------------------------ 2x(x^2-1) = 5x^2 + 4x -1 ------------ 2x(x^2-1) = (5x - 1 )(x +1) ----------- 2x(x+1)(x-1) = 5x-1 ----- 2x(x-1)

OpenStudy (4meisu):

Wouldn't you factorize 5x^2 + 4x - 1 as (5x-1) (x+5) though?